# Nilpotent structures of oriented neutral vector bundles

### v2

@@ 1 Introduction
A nilpotent structure $N$ gives a light-like subbundle $\pi_{N}$ of $(E,h)$ of rank $2n$. In addition, $N$ gives a section $\xi_{N}$ of $\bigwedge^{2n}\!E$, which is constructed by a local frame field of $\pi_{N}$ given by an admissible frame field of $N$. The fact that $\xi_{N}$ is well-defined is related to a Lie subgroup $G$ of $SO(2n,2n)$. The transition function between two admissible frame fields of $N$ is valued in $G$. It gives a local section of ${\rm End}\,\pi_{N}$ and its determinant is identically equal to one ([3]). If $n=1$, then the section of one of the light-like twistor spaces corresponding to $N$ is given by $(1/\sqrt{2})\xi_{N}$ ([2]).
@@
Let $\nabla$ be an $h$-connection of $E$, i.e., a connection of $E$ satisfying $\nabla h=0$. Then the Walker condition of $N$ is defined, based on the definition of Walker manifolds (see [6], [8], [9], [11], [19] for Walker manifolds). The Walker condition of $N$ is characterized by ^$\hat{\nabla}\xi_{N}=\alpha\otimes\xi_{N}$ with a $1$-form $\alpha$ on $M$ for the connection $\hat{\nabla}$ of $\bigwedge^{2n}\!E$ induced by $\nabla$ ((a) in Proposition 3.1). Since $\nabla N=0$ means ^$\hat{\nabla}\xi_{N}=0$ ((b) in Proposition 3.1), a nilpotent structure $N$ parallel with respect to $\nabla$ satisfies the Walker condition (see [11]). If $n=1$, then $\nabla N=0$ is equivalent to ^$\hat{\nabla}\xi_{N}=0$ ([2]). There exists a nilpotent structure $N$ satisfying the Walker condition and ^$\hat{\nabla}\xi_{N}\not=0$ (Example 3.1, Example 3.2, Example 3.3). In addition, if $n\geq 2$, then there exists a nilpotent structure $N$ satisfying ^$\hat{\nabla}\xi_{N}=0$ and $\nabla N\not=0$ (Example 3.4). Such an example $N$ in Example 3.1 $\sim$ Example 3.4 admits a nilpotent structure $N^{\prime}$ of $(E,h)$ satisfying the Walker condition and $E=\pi_{N}\oplus\pi_{N^{\prime}}$.
@@
In the present paper, we define special nilpotent structures related to a Lie subgroup $K$ of $SO(2n,2n)$. Each of them gives a principal $K$-bundle $P$ associated with $E$ and special admissible frame fields we use are given by local sections of $P$. In addition, $\nabla$ gives a connection in $P$, so that the connection form of $\nabla$ with respect to such a local section is valued in the Lie algebra of $K$. We call such a special nilpotent structure a $K$-nilpotent structure of $(E,h,\nabla)$. If $\nabla$ is flat and if $K\subset G$, then for a $K$-nilpotent structure of $(E,h,\nabla)$, we can find a local section $e$ of $P$ such that each local section of $E$ which appears in $e$ is parallel with respect to $\nabla$ (Proposition 4.2). A $G$-nilpotent structure of $(E,h,\nabla)$ is just a nilpotent structure of $(E,h)$ parallel with respect to $\nabla$ (Proposition 4.3). Let $H$ be a Lie subgroup of $G$ defined as in (4.1) below. If there exist a complex structure $I$ and paracomplex structures $J_{1}$, $J_{2}$ of $E$ such that $h$, $\nabla$, $I$, $J_{1}$, $J_{2}$ form a neutral hyperKähler structure of $E$, then $r(I-(\sin\theta)J_{1}+(\cos\theta)J_{2})$ ($r\in\mbox{\boldmath{$R$}}\setminus\{0\}$, $\theta\in[0,2\pi)$) are $H$-nilpotent structures of $(E,h,\nabla)$ (Theorem 4.4). See [5], [14] for paraquaternionic structures, and see [10], [17] for neutral hyperKähler $4$-manifolds. An $H$-nilpotent structure $N$ of $(E,h,\nabla)$ defines a unique light-like subbundle $\pi^{*}_{N}$ of $(E,h)$ of rank $2n$ satisfying $E=\pi_{N}\oplus\pi^{*}_{N}$, and a unique $H$-nilpotent structure $N^{*}$ of $(E,h,\nabla)$ (the dual $H$-nilpotent structure of $N$) satisfying $\pi_{N^{*}}=\pi^{*}_{N}$ (Theorem 4.5). Based on this, we observe that if there exists an $H$-nilpotent structure $N$ of $(E,h,\nabla)$, then $h$, $\nabla$, $I:=(1/2)(N+N^{*})$, $J_{1}:=-IJ_{2}$, $J_{2}:=(1/2)(N-N^{*})$ form a neutral hyperKähler structure of $E$ (Corollary 4.6).
@@ 2 Nilpotent structures
Let $N$ be an $\varepsilon$-nilpotent structure of $(E,h)$. Then such a frame field as $e$ is called an admissible frame field of $N$. Let $f=(f_{1},\dots,f_{2n},f_{2n+1},\dots,f_{4n})$ be an ordered pseudo-orthonormal local frame field of $(E,h)$ giving the orientation of $E$. Then $f$ is an admissible frame field of $N$ if and only if for each admissible frame field $e$ of $N$, an $SO(2n,2n)$-valued function $A$ on the intersection of the domains of $e$ and $f$ given by $fI^{\prime}_{4n,\varepsilon}=eI^{\prime}_{4n,\varepsilon}A$ is valued in the Lie subgroup $G$ of $SO(2n,2n)$ defined by $A_{0}\Lambda_{n}=\Lambda_{n}A_{0}$ for $A_{0}\in SO(2n,2n)$ ([3]). For an admissible frame field $e$ of $N$, we set
$\begin{array}[]{lcl}\xi_{1}:=e_{1}-e_{2n+1},&&\xi_{i}:=e_{i}-e_{2n+i},\\ \xi_{n+1}:=e_{n+1}+\varepsilon e_{3n+1},&&\xi_{n+i}:=e_{n+i}+e_{3n+i}\end{array}\ (i=2,\dots,n)$ (2.1)
and $\xi=\xi_{1}\wedge\dots\wedge\xi_{2n}$. Then $\xi$ does not depend on the choice of an admissible frame field $e$ of $N$ ([3]). Therefore $N$ gives a section $\xi_{N}$ of the $2n$-fold exterior power $\bigwedge^{2n}\!E$ of $E$. A section $N$ of ${\rm End}\,E$ is a nilpotent structure of $(E,h)$ if and only if $N$ satisfies
@@
([2], [3]). In particular, $N$ gives a null structure on each fiber of $E$ and $h$ is null-Hermitian with respect to $N$ (see [11]). The section $\xi_{N}$ of $\bigwedge^{2n}\!E$ given by $N$ corresponds to the subbundle $\pi_{N}$ with the orientation assigned by $\xi_{N}$.
@@ 3 The Walker condition
We denote by $\hat{\nabla}$ the connection of $\bigwedge^{2n}\!E$ induced by $\nabla$. We will prove
@@
    if $N$ is parallel with respect to $\nabla$, then ^$\hat{\nabla}\xi_{N}=0$, and therefore $N$ satisfies the Walker condition with respect to $\nabla$.
@@
If $N$ satisfies ^$\hat{\nabla}\xi_{N}=\alpha\otimes\xi_{N}$ for a $1$-form $\alpha$ on $M$, then each $\nabla\xi_{i}$ is a $1$-form valued in $\pi_{N}$, which means that $N$ satisfies the Walker condition. Hence we obtain (a) in Proposition 3.1. Suppose that $N$ satisfies $\nabla N=0$. Then we have $\omega\Lambda_{n}=\Lambda_{n}\omega$. This is rewritten into
@@
Therefore we obtain
$\begin{split}&(\nabla\xi_{1}\ \dots\ \nabla\xi_{n})\\ &=(\xi_{1}\ \dots\ \xi_{n})(D_{11}-D_{13})+(\xi_{n+1}\ \dots\ \xi_{2n})(D_{21}-D_{23}),\\ &(\nabla\xi_{n+1}\ \dots\ \nabla\xi_{2n})\\ &=(\xi_{1}\ \dots\ \xi_{n})(D_{12}+D_{14})+(\xi_{n+1}\ \dots\ \xi_{2n})(D_{22}+D_{24}).\end{split}$ (3.5)
Since
$D_{22}+D_{24}=-{}^{t}(D_{22}-D_{42})=-{}^{t}(D_{11}-D_{13})$
by (3.1) and (3.4), the sum of the $(i,i)$-components of $D_{22}+D_{24}$ and $D_{11}-D_{13}$ is zero for $i=1,\dots,n$. Therefore ^$\hat{\nabla}\xi_{N}=0$. Hence we obtain (b) in Proposition 3.1. $\square$
@@
In particular, if $dF=(1/2)dfI_{n}$ for a function $f$, then $\nabla N=df\otimes J$, where $J$ is a paracomplex structure of $E$ defined by
$JeI^{\prime}_{4n,\varepsilon}=eI^{\prime}_{4n,\varepsilon}\left[\begin{array}[]{cccc}O_{n}&O_{n}&I_{n}&O_{n}\\ O_{n}&O_{n}&O_{n}&-I_{n}\\ I_{n}&O_{n}&O_{n}&O_{n}\\ O_{n}&-I_{n}&O_{n}&O_{n}\end{array}\right].$
In addition, if we suppose $n=1$, then $\nabla N=df\otimes J$ corresponds to the relation ^$\hat{\nabla}\Omega_{N}=df\otimes\Omega_{J}$, where $\Omega_{N}$ is a section of $U_{0}\!\left(\bigwedge^{2}_{\varepsilon}\!E\right)$ corresponding to $N$ and given by $(1/\sqrt{2})\xi_{N}$, and $\Omega_{J}$ is a section of $U_{-}\!\left(\bigwedge^{2}_{\varepsilon}\!E\right)$ corresponding to $J$.

### v3

@@ 1 Introduction
A nilpotent structure $N$ gives a light-like subbundle $\pi_{N}$ of $(E,h)$ of rank $2n$. In addition, $N$ gives a nowhere zero section $\xi_{N}$ of a line bundle $\bigwedge^{2n}\!\pi_{N}$, which is constructed by a local frame field of $\pi_{N}$ given by an admissible frame field of $N$. The fact that $\xi_{N}$ is well-defined is related to a Lie subgroup $G$ of $SO(2n,2n)$. The transition function between two admissible frame fields of $N$ is valued in $G$. It gives a local section of ${\rm End}\,\pi_{N}$ and its determinant is identically equal to one ([4]). In the present paper, we will see that $N$ gives a nondegenerate section $\Theta_{N}$ of $\bigwedge^{2}\!\pi_{N}$ satisfying $\xi_{N}=((-1)^{\frac{n(n-1)}{2}}/n!)\Theta^{n}_{N}$ and that a nondegenerate section $\Theta$ of the $2$-fold exterior power of a light-like subbundle $L$ of $(E,h)$ of rank $2n$ gives a nilpotent structure $N$ of $(E,h)$ satisfying $\pi_{N}=L$ and $\Theta_{N}=\Theta$ (Theorem 2.1). If $n=1$, then the section of one of the light-like twistor spaces corresponding to $N$ is given by $(1/\sqrt{2})\xi_{N}$ ([3]).
@@
Let $\nabla$ be an $h$-connection of $E$, i.e., a connection of $E$ satisfying $\nabla h=0$. Then the Walker condition of $N$ is defined, based on the definition of Walker manifolds (see [8], [11], [12], [15], [30] for Walker manifolds). The Walker condition of $N$ is characterized by ^$\hat{\nabla}\xi_{N}=\alpha\otimes\xi_{N}$ with a $1$-form $\alpha$ on $M$ for the connection $\hat{\nabla}$ of $\bigwedge^{2n}\!E$ induced by $\nabla$ ((a) in Proposition 3.1). Therefore the Walker condition just means that $\hat{\nabla}$ induces a connection of $\bigwedge^{2n}\!\pi_{N}$. If we also denote by $\hat{\nabla}$ the induced connection of $\bigwedge^{2}\!E$, then $\nabla N=0$ is equivalent to ^$\hat{\nabla}\Theta_{N}=0$, and therefore $\nabla N=0$ yields ^$\hat{\nabla}\xi_{N}=0$ ((b) in Proposition 3.1), which means that a nilpotent structure $N$ parallel with respect to $\nabla$ satisfies the Walker condition (see [15]). If $n=1$, then $\nabla N=0$ is equivalent to ^$\hat{\nabla}\xi_{N}=0$ ([3]). There exists a nilpotent structure $N$ satisfying the Walker condition and ^$\hat{\nabla}\xi_{N}\not=0$ (Example 3.1, Example 3.2, Example 3.3). In addition, if $n\geq 2$, then there exists a nilpotent structure $N$ satisfying ^$\hat{\nabla}\xi_{N}=0$ and $\nabla N\not=0$ (Example 3.4). Such an example $N$ in Example 3.1 $\sim$ Example 3.4 admits a nilpotent structure $N^{\prime}$ of $(E,h)$ satisfying the Walker condition and $E=\pi_{N}\oplus\pi_{N^{\prime}}$.
@@
In the present paper, we define special nilpotent structures related to a Lie subgroup $K$ of $SO(2n,2n)$. Each of them gives a principal $K$-bundle $P$ associated with $E$ and special admissible frame fields we use are given by local sections of $P$. In addition, $\nabla$ gives a connection in $P$, so that the connection form of $\nabla$ with respect to such a local section is valued in the Lie algebra of $K$. We call such a special nilpotent structure a $K$-nilpotent structure of $(E,h,\nabla)$. If $\nabla$ is flat and if $K\subset G$, then for a $K$-nilpotent structure of $(E,h,\nabla)$, we can find a local section $e$ of $P$ such that each local section of $E$ which appears in $e$ is parallel with respect to $\nabla$ (Proposition 4.2). A $G$-nilpotent structure of $(E,h,\nabla)$ is just a nilpotent structure of $(E,h)$ parallel with respect to $\nabla$ (Proposition 4.3). Let $H$ be a Lie subgroup of $G$ defined as in (4.1) below. If there exist a complex structure $I$ and paracomplex structures $J_{1}$, $J_{2}$ of $E$ such that $h$, $\nabla$, $I$, $J_{1}$, $J_{2}$ form a neutral hyperKähler structure of $E$, then $r(I-(\sin\theta)J_{1}+(\cos\theta)J_{2})$ ($r\in\mbox{\boldmath{$R$}}\setminus\{0\}$, $\theta\in[0,2\pi)$) are $H$-nilpotent structures of $(E,h,\nabla)$ (Theorem 4.4). See [7], [21] for paraquaternionic structures, and see [13], [25] for neutral hyperKähler $4$-manifolds. An $H$-nilpotent structure $N$ of $(E,h,\nabla)$ defines a unique light-like subbundle $\pi^{\times}_{N}$ of $(E,h)$ of rank $2n$ satisfying $E=\pi_{N}\oplus\pi^{\times}_{N}$, and a unique $H$-nilpotent structure $N^{\times}$ of $(E,h,\nabla)$ (the dual $H$-nilpotent structure of $N$) satisfying $\pi_{N^{\times}}=\pi^{\times}_{N}$ (Theorem 4.5). Based on this, we observe that if there exists an $H$-nilpotent structure $N$ of $(E,h,\nabla)$, then $h$, $\nabla$, $I:=(1/2)(N+N^{\times})$, $J_{1}:=-IJ_{2}$, $J_{2}:=(1/2)(N-N^{\times})$ form a neutral hyperKähler structure of $E$ (Corollary 4.6).
@@
Remark 1.2
The metrics of vector bundles in the present paper are neutral. We can refer to [18] for almost paracomplex structures on Riemannian or neutral 4-manifolds. See [14] for a characterization of anti-self-dual null-Kähler $4$-manifolds. Neutral metrics appear in the studies of spaces of oriented geodesics ([1], [20], [28], [29]), the ultra-hyperbolic equation ([6], [10], [19], [24]) and quantum field theories ([27]).
@@ 2 Nilpotent structures
Let $N$ be an $\varepsilon$-nilpotent structure of $(E,h)$. Then such a frame field as $e$ is called an admissible frame field of $N$. Let $f=(f_{1},\dots,f_{2n},f_{2n+1},\dots,f_{4n})$ be an ordered pseudo-orthonormal local frame field of $(E,h)$ giving the orientation of $E$. Then $f$ is an admissible frame field of $N$ if and only if for each admissible frame field $e$ of $N$, an $SO(2n,2n)$-valued function $A$ on the intersection of the domains of $e$ and $f$ given by $fI^{\prime}_{4n,\varepsilon}=eI^{\prime}_{4n,\varepsilon}A$ is valued in the Lie subgroup $G$ of $SO(2n,2n)$ defined by $A_{0}\Lambda_{n}=\Lambda_{n}A_{0}$ for $A_{0}\in SO(2n,2n)$ ([4]). A section $N$ of ${\rm End}\,E$ is a nilpotent structure of $(E,h)$ if and only if $N$ satisfies
@@
([3], [4]). In particular, $N$ gives a null structure on each fiber of $E$ and $h$ is null-Hermitian with respect to $N$ (see [15]).
Let $E^{*}$ be the dual bundle of $E$. Then $h$ induces a neutral metric $h^{*}$ of $E^{*}$. For an ordered pseudo-orthonormal local frame field $e=(e_{1},\dots,e_{2n},e_{2n+1},\dots,e_{4n})$ of $(E,h)$, let $e^{*}:=(e^{1},\dots,e^{2n},e^{2n+1},\dots,e^{4n})$ be the dual frame field of $e$. Then $e^{*}$ is an ordered pseudo-orthonormal local frame field of $(E^{*},h^{*})$. Let $N$ be an $\varepsilon$-nilpotent structure of $(E,h)$. Then $N$ gives a section $N^{*}$ of ${\rm End}\,E^{*}$ by
$N^{*}\phi^{*}:=\phi^{*}\circ N$ (2.1)
for a local section $\phi^{*}$ of $E^{*}$. Let $e$ be an admissible frame field of $N$. Then $e^{*}$ satisfies $N^{*}e^{*}I^{\prime}_{4n,\varepsilon}=e^{*}I^{\prime}_{4n,\varepsilon}{}^{t}\!\Lambda_{n}$. By this, we obtain $h^{*}(\phi^{*},N^{*}\phi^{*})=0$ for any local section $\phi^{*}$ of $E^{*}$. Therefore, if we set
$\Theta_{N}(\phi^{*},\psi^{*}):=h^{*}(\phi^{*},N^{*}\psi^{*})$ (2.2)
for local sections $\phi^{*}$, $\psi^{*}$ of $E^{*}$, then $\Theta_{N}(\psi^{*},\phi^{*})=-\Theta_{N}(\phi^{*},\psi^{*})$. This means that $\Theta_{N}$ is a section of the $2$-fold exterior power $\bigwedge^{2}\!E$ of $E$ given by $N$. In addition, by $N^{*}e^{*}I^{\prime}_{4n,\varepsilon}=e^{*}I^{\prime}_{4n,\varepsilon}{}^{t}\!\Lambda_{n}$, we obtain
$\Theta_{N}=\sum^{n}_{i=1}\xi_{i}\wedge\xi_{n+i},$
where
$\begin{array}[]{lcl}\xi_{1}:=e_{1}-e_{2n+1},&&\xi_{i}:=e_{i}-e_{2n+i},\\ \xi_{n+1}:=e_{n+1}+\varepsilon e_{3n+1},&&\xi_{n+i}:=e_{n+i}+e_{3n+i}\end{array}\ (i=2,\dots,n).$ (2.3)
Therefore $\Theta_{N}$ is a section of $\bigwedge^{2}\!\pi_{N}$. In addition, $\Theta_{N}$ is nondegenerate, that is, $\Theta_{N}$ gives a nowhere zero section $\xi_{N}$ of $\bigwedge^{2n}\!\pi_{N}$ by
$\xi_{N}=\dfrac{(-1)^{\frac{n(n-1)}{2}}}{n!}\Theta^{n}_{N}=\xi_{1}\wedge\dots\wedge\xi_{2n}.$ (2.4)
Let $L$ be a subbundle of $E$ of rank $2n$. We say that $L$ is an $\varepsilon$-light-like subbundle of $(E,h)$ if on a neighborhood of each point of $M$, there exists an ordered pseudo-orthonormal local frame field $e=(e_{1},\dots,e_{2n},e_{2n+1},\dots,e_{4n})$ of $(E,h)$ giving the orientation of $E$ such that $\xi_{1},\dots,\xi_{n},\xi_{n+1},\dots,\xi_{2n}$ as in (2.3) form a local frame field of $L$. Let $L$ be an $\varepsilon$-light-like subbundle of $(E,h)$. Let $\Theta$ be a nondegenerate section of $\bigwedge^{2}\!L$. Then on a neighborhood of each point of $M$, there exists an ordered pseudo-orthonormal local frame field $e$ of $(E,h)$ giving the orientation of $E$ and satisfying $\Theta=\sum^{n}_{i=1}\xi_{i}\wedge\xi_{n+i}$ with (2.3). For $\Theta$, let $N^{*}$ be a section of ${\rm End}\,E^{*}$ given by $\Theta(\phi^{*},\psi^{*})=h^{*}(\phi^{*},N^{*}\psi^{*})$ for local sections $\phi^{*}$, $\psi^{*}$ of $E^{*}$. Then $N^{*}$ satisfies $h^{*}(\phi^{*},N^{*}\phi^{*})=0$ and $N^{*}e^{*}I^{\prime}_{4n,\varepsilon}=e^{*}I^{\prime}_{4n,\varepsilon}{}^{t}\!\Lambda_{n}$. For $N^{*}$, let $N$ be a section of ${\rm End}\,E$ given by $N^{*}\phi^{*}=\phi^{*}\circ N$. Then $N$ is an $\varepsilon$-nilpotent structure of $(E,h)$ such that $e$ is an admissible frame field of $N$ and $N$ satisfies $\pi_{N}=L$ and $\Theta_{N}=\Theta$. Hence we obtain
Theorem 2.1
There exists a one-to-one correspondence between the set of $\varepsilon$-nilpotent structures of $(E,h)$ and the set of nondegenerate sections of the $2$-fold exterior powers of $\varepsilon$-light-like subbundles of $(E,h)$:::: each $\varepsilon$-nilpotent structure $N$ corresponds to a nondegenerate section $\Theta_{N}$ of $\bigwedge^{2}\!\pi_{N}$ by (2.1) and (2.2).
Remark 2.1
An $\varepsilon$-nilpotent structure $N$ of $(E,h)$ gives a nondegenerate section $\Theta_{N}$ of $\bigwedge^{2}\!\pi_{N}$. Therefore $\xi_{N}$ in (2.4) does not depend on the choice of an admissible frame field $e$ of $N$, which was already obtained in [4].
@@ 3 The Walker condition
The connection $\nabla$ induces connections of $\bigwedge^{2}\!E$ and $\bigwedge^{2n}\!E$, which are denoted by $\hat{\nabla}$. We will prove
@@
    $N$ is parallel with respect to $\nabla$ if and only if the corresponding section $\Theta_{N}$ of $\bigwedge^{2}\!\pi_{N}$ is horizontal with respect to $\hat{\nabla}$, and if we suppose one of these conditions, then ^$\hat{\nabla}\xi_{N}=0$, and therefore $N$ satisfies the Walker condition with respect to $\nabla$.
@@
If $N$ satisfies ^$\hat{\nabla}\xi_{N}=\alpha\otimes\xi_{N}$ for a $1$-form $\alpha$ on $M$, then each $\nabla\xi_{i}$ is a $1$-form valued in $\pi_{N}$, which means that $N$ satisfies the Walker condition. Hence we obtain (a) in Proposition 3.1. The condition $\nabla N=0$ is equivalent to $\omega\Lambda_{n}=\Lambda_{n}\omega$. This can be rewritten as
@@
We see that (3.4) is equivalent to ^$\hat{\nabla}\Theta_{N}=0$. Therefore $\nabla N=0$ is equivalent to ^$\hat{\nabla}\Theta_{N}=0$. If we suppose ^$\hat{\nabla}\Theta_{N}=0$, then by (2.4), we obtain ^$\hat{\nabla}\xi_{N}=0$. Hence we obtain (b) in Proposition 3.1. $\square$
@@
Remark 3.4
Let $E$, $h$ be as in the beginning of Section 2. Let $J$ be a section of ${\rm End}\,E$. As in [4], we say that $J$ is an $\varepsilon$-paracomplex structure of $(E,h)$ if $J$ satisfies
  • (i)
    $J$ is a paracomplex structure of $E$,
  • (ii)
    $J$ is $h$-reversing, that is, $J^{*}h=-h$,
  • (iii)
    on a neighborhood of each point of $M$, there exists an ordered pseudo-orthonormal local frame field $e=(e_{1},\dots e_{2n},e_{2n+1},\dots,e_{4n})$ of $E$ giving the orientation and satisfying
    $JeI^{\prime}_{4n,\varepsilon}=eI^{\prime}_{4n,\varepsilon}\Lambda_{n,-},\quad\Lambda_{n,-}:=\left[\begin{array}[]{cccc}O_{n}&O_{n}&I_{n}&O_{n}\\ O_{n}&O_{n}&O_{n}&-I_{n}\\ I_{n}&O_{n}&O_{n}&O_{n}\\ O_{n}&-I_{n}&O_{n}&O_{n}\end{array}\right].$ (3.6)
Let $J$ be an $\varepsilon$-paracomplex structure of $(E,h)$. Then such a frame field as $e$ is called an admissible frame field of $J$. We see that $J$ gives a section $J^{*}$ of ${\rm End}\,E^{*}$ by $J^{*}\phi^{*}=\phi^{*}\circ J$ for a local section $\phi^{*}$ of $E^{*}$. Then we have $J^{*}e^{*}I^{\prime}_{4n,\varepsilon}=e^{*}I^{\prime}_{4n,\varepsilon}\Lambda_{n,-}$. Therefore, if we set $\Theta_{J}(\phi^{*},\psi^{*}):=h^{*}(\phi^{*},J^{*}\psi^{*})$, then $\Theta_{J}$ is a section of $\bigwedge^{2}\!E$ and locally represented as
$\Theta_{J}=\sum^{n}_{i=1}(e_{i}\wedge e_{2n+i}+(\varepsilon 1)^{\delta_{i1}}e_{3n+i}\wedge e_{n+i}).$ (3.7)
Let $\nabla$ be an $h$-connection of $E$. Then $\nabla J=0$ is equivalent to ^$\hat{\nabla}\Theta_{J}=0$. Suppose that $\nabla J$ is locally represented as the tensor product of a $1$-form $\alpha$ and an $\varepsilon$-nilpotent structure $N$ so that $e$ is an admissible frame field of both $J$ and $N$. Then
$\begin{split}&\omega^{n+i}_{j}+(\varepsilon 1)^{\delta_{i1}}\omega^{3n+i}_{2n+j}=\omega^{n+i}_{2n+j}+(\varepsilon 1)^{\delta_{i1}}\omega^{3n+i}_{j}=\alpha\delta^{i}_{j},\\ &\omega^{i}_{j}-\omega^{2n+i}_{2n+j}=\omega^{i}_{2n+j}-\omega^{2n+i}_{j}=0,\\ &\omega^{n+i}_{n+j}-(\varepsilon 1)^{\delta_{i1}+\delta_{j1}}\omega^{3n+i}_{3n+j}=\omega^{n+i}_{3n+j}-(\varepsilon 1)^{\delta_{i1}+\delta_{j1}}\omega^{3n+i}_{n+j}=0\end{split}$ (3.8)
are obtained in [4] for $i,j=1,\dots,n$. Therefore by (3.7) and (3.8), we obtain ^$\hat{\nabla}\Theta_{J}=\alpha\otimes\Theta_{N}$.
@@
In particular, if $dF=(1/2)dfI_{n}$ for a function $f$, then $\nabla N=df\otimes J$, where $J$ is an $\varepsilon$-paracomplex structure of $(E,h)$ defined by (3.6), and ^$\hat{\nabla}\Theta_{N}=df\otimes\Theta_{J}$. Therefore, if we suppose $n=1$, then ^$\hat{\nabla}\Omega_{N}=df\otimes\Omega_{J}$, where $\Omega_{N}$ is the section of $U_{0}\!\left(\bigwedge^{2}_{\varepsilon}\!E\right)$ corresponding to $N$ and given by $(1/\sqrt{2})\xi_{N}$, and $\Omega_{J}$ is the section of $U_{-}\!\left(\bigwedge^{2}_{\varepsilon}\!E\right)$ corresponding to $J$.